An evaporating dish weighs 44.3170 grams. After 8.5 mL of salt water was added to the dish, the combination of the dish and salt water weighed 52.987 grams. The dish was then heated until a constant mass of 44.5990 grams was obtained. What was the mass percent of salt dissolved in the salt water


1
Expert's answer
2021-12-29T11:03:01-0500

Mass of salt water "=52.987-44.317"

"=8.67g"


Mass of salt "=44.5990-44.317=0\n.282g"


Mass percentage "=\frac{0.282}{8.67}×100\%"

"=3.2526\%"